OpenStudy (anonymous):

How do you find the range of the function 4x+3y= −10 with a domain of x> −1?

jimthompson5910 (jim_thompson5910):

First solve for y 4x+3y= -10 3y= -4x-10 y= -4/3x-10/3 So this is a linear equation where it slopes downward (as you read it from left to right)

jimthompson5910 (jim_thompson5910):

To find the range, you plug in x = -1 and evaluate to find y. This will give you the largest value in the range. The range will be the set of numbers smaller than this evaluated y value.

OpenStudy (anonymous):

So the range would be y< 4/3?

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

Still a bit confused- what do I do after I plug in the x value? And why did I plug in the x value?

jimthompson5910 (jim_thompson5910):

x = -1 is the endpoint of the domain

jimthompson5910 (jim_thompson5910):

so plugging that into the equation will give you the endpoint of the range

OpenStudy (anonymous):

Okay I get that, so after you plug it in you get y = 4/3 − 10/3 so what comes next?

jimthompson5910 (jim_thompson5910):

y = 4/3 − 10/3 y = (4-10)/3 y = ???

OpenStudy (anonymous):

y is 2??

jimthompson5910 (jim_thompson5910):

close, y is -2

OpenStudy (anonymous):

Oh sorry i missed the negative sign but THANK YOU SO MUCH! I have a quiz tomorrow and I was confused!!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Hey just one more thing (sorry!)- How do I know whether it is < or >, like after i figure out the number value of y?